OpenStudy (hba):

Determine the following functions as even,odd or neither. (a) f(x)=|x| (b) f(x)=x+1 (c) f(x)=x^5-x/1+x

hartnn (hartnn):

if f(x) = f(-x) then function is even

OpenStudy (hba):

Answers: (a)even (b)and (c)Neither

Parth (parthkohli):

\[f(x) =f(-x) \]Even^^\[f(x) = -f(x)\]Odd^

OpenStudy (hba):

Right or wrong?

hartnn (hartnn):

if f(-x)=-f(x) then function is odd

OpenStudy (hba):

I know all that -_-

OpenStudy (hba):

Just check my answers

hartnn (hartnn):

whats and ?

Parth (parthkohli):

Abs value function has \(f(x) = f(-x) \ \ \checkmark \)

hartnn (hartnn):

a and c are correct what about b?

Parth (parthkohli):

For second, \(f(x) \ne f(-x)\) and \(f(x) \ne - f(x)\) for all \(x\). What do you think?

OpenStudy (hba):

f(-x)=-x+1 so,Neither even nor odd

hartnn (hartnn):

correct.

Parth (parthkohli):

Yeah!

OpenStudy (hba):

Thanks for the help guys :)